# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ B.1. Does dispositional trust affect decision-making and fairness perception measures?
As discussed in section 5.2, we include a measurement of a participant’s dispositional trust in AI as a fixed effect in our linear models. The effects and their significance was generally not consistent across models (See Table 2). Metric Bias Type Model $F$ $p$ Coef Std Error Expl $F(1,115)=4.07$ $0.0459$ $1.31$ $0.592$ Disclosure $F(1,49)=5.86$ $0.0192$ $1.70$ $0.700$ protected Disclosure with Explanation $F(1,66)=1.6$ $0.2104$ $0.89$ $0.702$ Expl $F(1,227)=32.7$ $0.0000$** $1.54$ $0.280$ Disclosure $F(1,91)=25.8$ $0.0000$** $1.69$ $0.332$ Fairness Rating proxy Disclosure with Explanation $F(1,136)=10.11$ $0.0018$** $1.13$ $0.355$ Expl $F(1,115)=0.87$ $0.3540$ $-0.31$ $0.284$ Disclosure $F(1,49)=0.18$ $0.6737$ $-0.15$ $0.354$ protected Disclosure with Explanation $F(1,66)=0.85$ $0.3598$ $-0.30$ $0.331$ Expl $F(1,227)=2.52$ $0.1135$ $-0.12$ $0.087$ Disclosure $F(1,91)=9.11$ $0.0033$** $-0.45$ $0.150$ Fairness Saliency proxy Disclosure with Explanation $F(1,136)=0.01$ $0.9361$ $-0.01$ $0.125$ Expl $F(1,115)=6.25$ $0.0138$** $-0.35$ $0.149$ Disclosure $F(1,99)=0.53$ $0.4681$ $-0.15$ $0.208$ protected Disclosure with Explanation $F(1,133)=3.27$ $0.0730$ $-0.28$ $0.153$ Expl $F(1,227)=5.73$ $0.0175$** $-0.22$ $0.112$ Disclosure $F(1,182)=0.31$ $0.5778$ $0.07$ $0.128$ Parity proxy Disclosure with Explanation $F(1,135)=3.55$ $0.0618$ $-0.19$ $0.100$ Expl $F(1,115)=2.09$ $0.1511$ $0.04$ $0.031$ Disclosure $F(1,49)=4.71$ $0.0349$ $0.10$ $0.044$ protected Disclosure with Explanation $F(1,66)=0.23$ $0.6366$ $0.02$ $0.033$ Expl $F(1,227)=2.97$ $0.0863$ $0.02$ $0.018$ Disclosure $F(1,91)=1.73$ $0.1917$ $0.03$ $0.026$ Accuracy proxy Disclosure with Explanation $F(1,136)=2.22$ $0.1382$ $0.03$ $0.019$ Expl $F(1,115)=0.06$ $0.8126$ $0.01$ $0.096$ Disclosure $F(1,49)=0.01$ $0.9113$ $-0.01$ $0.125$ protected Disclosure with Explanation $F(1,66)=0.53$ $0.4674$ $0.08$ $0.104$ Expl $F(1,227)=8.8$ $0.0033$** $0.14$ $0.054$ Disclosure $F(1,91)=0.6$ $0.4400$ $0.06$ $0.072$ FNR proxy Disclosure with Explanation $F(1,136)=9.69$ $0.0023$** $0.20$ $0.063$ Expl $F(1,115)=0.7$ $0.4051$ $-0.08$ $0.107$ Disclosure $F(1,49)=0.73$ $0.3983$ $-0.12$ $0.144$ protected Disclosure with Explanation $F(1,66)=0.59$ $0.4458$ $-0.10$ $0.125$ Expl $F(1,227)=9.44$ $0.0024$** $-0.16$ $0.059$ Disclosure $F(1,91)=2.07$ $0.1533$ $-0.12$ $0.085$ FPR proxy Disclosure with Explanation $F(1,136)=8.79$ $0.0036$** $-0.22$ $0.073$ Table 2. Effects of dispositional trust in AI on different outcome metrics (Perception, Parity, Accuracy, FPR, and FNR). We consistently find that dispositional trust has no significant impact on fairness ratings in the protected conditions, while it significantly increases fairness ratings in the proxy conditions. In other words, when biases are direct, people are equally able to notice model biases even when they tend to trust AI in general; however, when biases are indirect, people with higher dispositional trust in AI are less likely to believe that the model is unfair. For participants’ fairness saliency, we see that there is never a significant effect in the case of direct bias, but higher dispositional trust significantly decreased the rate of disagreement only when considering disclosures without explanations. We also find that, under models that consider the effect of explanations and disclosures with explanations, increases dispositional trust in AI significantly increased FPR and decreased FNR in the proxy conditions only. This is likely due to participants with higher trust in AI being more influenced by subtle indirect biases leading to lower acceptance rates for female applicants. This is also supported by the models measuring the effect of explanations on gender parity. Here, we see that an increased dispositional trust in AI significantly decreased parity under both types of bias. Overall, people with a greater dispositional trust in AI tended to make more unfair decisions (when working with a biased model) and were less likely to notice indirect bias.
@@ B.2. Do explanations and disclosures affect learned trust?
In this section, we discuss the effect of our interventions—explanations, disclosures without explanations, and disclosures with explanations (§5.1.3)—on learned trust over dispositional trust in AI generally. We perform statistical tests similar to the ones described in §5.2. We consider the different trust measures as the dependent variable and the treatment as the fixed effect term. We also control for the dispositional trust level as a fixed effect. As seen in Table 6, we find that our treatments generally have no effect on trust ratings in models exhibiting direct bias, except for explanations alone leading to significantly lowered feelings that the AI system works well. In the case of indirect bias, we often see that full disclosure with explanations (and sometimes also full disclosure without explanations or bias disclosure with explanations) has a significant effect on learned trust. These effects demonstrate lowered feelings that the AI system works well as well as decreased feelings that the AI system can perform as well as an untrained human, decreased confidence in the system, decreased feelings of safety when relying on the system, and increased wariness of the AI system. We also find that explanations alone significantly increase participant’s perception of model predictability in the proxy conditions but not the protected conditions. Without explanations, full bias and correlation disclosure also significantly increased predictability. Likely due to explanations increasing predictability on its own, no disclosure significantly affected predictability when paired with explanations. This is to say that our models are already seen as relatively predictable when biases are direct, but when biases are indirect, explanations or disclosure of model bias and the model’s usage of the university feature help make the model more predictable. better confidence predictable safe wary works well Feature Phase Expl? Disclosure sig coef sig coef sig coef sig coef sig coef sig coef $1$** $-0.61$** $-0.43$** $-0.51$** $-0.29$ $0.04$** $-0.43$ $1$** $-0.93$** $-0.66$** $-0.37$ $-0.26$ $0.10$** $-0.81$ $2$ Bias *** $-0.33$** $-0.23$** $-0.25$** $-0.20$ $0.02$** $-0.28$ Protected $2$ Bias *** $-0.51$** $-0.38$** $-0.19$** $-0.22$ $0.12$** $-0.48$ $1$** $-0.55$** $-0.34$** $-0.68$** $-0.22$ $-0.15$** $-0.32$ $1$** $-0.54$ $-0.19$ $0.01$ $0.03$ $-0.14$** $-0.29$ $2$ Bias *** $-0.39$** $-0.21$** $-0.29$ $-0.05$ $0.02$** $-0.25$ $2$ Full *** $-0.31$** $-0.28$** $-0.23$ $-0.13$ $0.01$** $-0.37$ $2$ Bias *** $-0.30$** $-0.29$ $0.02$ $-0.09$ $0.06$** $-0.30$ Proxy $2$ Full *** $-0.36$** $-0.20$ $-0.04$ $-0.08$ $0.00$** $-0.26$ Table 3. Comparison of trust in AI generally vs trust in our models in varied conditions and phases.
@@ B.3. Does dispositional trust differ from learned trust?
In the previous section, we discussed how interventions affected learned trust when controlling for baseline dispositional trust. Here, we study whether there is a significant difference between dispositional trust and learned trust in the biased models across the questions described in §4.2. These results are shown in Table 3. We find that participants usually thought our model worked worse than AI does generally, that it inspired less confidence, and was less predictable. Participants also regarded our AI system as less safe than AI in general, but this is primarily true only in the case of direct bias. Surprisingly, participants did not consider our AI systems to be less safe than general in phase 1 when they were given explanations (which would have directly indicated that the system used gender as a feature to determine loan outcomes). Participants’ wariness of the biased models was not significantly different from their baseline wariness in AI.

### v2

@@ A.2. Expected Accuracy, FNR, and FPR
Using our calculated probability of the ground-truth acceptance of a given applicant (with original features $\boldsymbol{x}$ and synthetic feature $x^{\ast}$), we can calculate an expected accuracy, FNR, and FPR for a given human-AI team. For example, for expected FNR, we consider the expected number of false negatives over the expected number of ground truth positives.
Expected FNR $=\frac{\mathbb{E}[\hbox{\# False Negatives}]}{\mathbb{E}[\hbox{\#Positives}]}$
@@
Expected $\text{Expected FNR}=\frac{\sum_{(i,j)\in S}P(Y_{i}=1|\boldsymbol{x}_{i},x^{\ast}_{i})\times(1-\hat{y}_{i,j})}{\sum_{i}P(Y_{i}=1|\boldsymbol{x}_{i},x^{\ast}_{i})}.$
Similarly, we can calculate the expected FPR and Accuracy.
Expected FPR $=\frac{\mathbb{E}[\hbox{\# False Positives}]}{\mathbb{E}[\hbox{\#Negatives}]}$
$=\frac{\sum_{(i,j)\in S}P(Y_{i}=0|\boldsymbol{x}_{i},x^{\ast}_{i})\times\hat{y}_{i,j}}{\sum_{i}P(Y_{i}=0|\boldsymbol{x}_{i},x^{\ast}_{i})}$
Expected Accuracy $=\frac{\mathbb{E}[\hbox{\# True Positives + \# True Negatives}]}{\mathbb{E}[\hbox{\# Samples}]}$
$=\frac{1}{\mid S\mid}\sum_{(i,j)\in S}\mathbb{E}[TP_{i,j}+TN_{i,j}]$
$=\frac{1}{\mid S\mid}\sum_{(i,j)\in S}\big{(}P(Y_{i}=1|\boldsymbol{x}_{i},x^{\ast}_{i})\times\hat{y}_{i,j}\big{)}$
$\quad+\big{(}P(Y_{i}=0|\boldsymbol{x}_{i},x^{\ast}_{i})\times(1-\hat{y}_{i,j})\big{)}$
@@ B.1. Does dispositional trust affect decision-making and fairness perception measures?
As discussed in section 5.2, we include a measurement of a participant’s dispositional trust in AI as a fixed effect in our linear models. The effects and their significance was generally not consistent across models (See Table 2).
Metric Feature Model $F$ $p$ Coef Std Error
Expl $F(1,115)=4.07$ $0.0459$ $1.31$ $0.592$
Disclosure $F(1,49)=5.86$ $0.0192$ $1.70$ $0.700$
protected Disclosure with Explanation $F(1,66)=1.6$ $0.2104$ $0.89$ $0.702$
Expl $F(1,227)=32.7$ $0.0000$** $1.54$ $0.280$
Disclosure $F(1,91)=25.8$ $0.0000$** $1.69$ $0.332$
Fairness Rating proxy Disclosure with Explanation $F(1,136)=10.11$ $0.0018$** $1.13$ $0.355$
Expl $F(1,115)=0.87$ $0.3540$ $-0.31$ $0.284$
Disclosure $F(1,49)=0.18$ $0.6737$ $-0.15$ $0.354$
protected Disclosure with Explanation $F(1,66)=0.85$ $0.3598$ $-0.30$ $0.331$
Expl $F(1,227)=2.52$ $0.1135$ $-0.12$ $0.087$
Disclosure $F(1,91)=9.11$ $0.0033$** $-0.45$ $0.150$
Fairness Saliency proxy Disclosure with Explanation $F(1,136)=0.01$ $0.9361$ $-0.01$ $0.125$
Expl $F(1,115)=6.25$ $0.0138$** $-0.35$ $0.149$
Disclosure $F(1,99)=0.53$ $0.4681$ $-0.15$ $0.208$
protected Disclosure with Explanation $F(1,133)=3.27$ $0.0730$ $-0.28$ $0.153$
Expl $F(1,227)=5.73$ $0.0175$** $-0.22$ $0.112$
Disclosure $F(1,182)=0.31$ $0.5778$ $0.07$ $0.128$
Parity proxy Disclosure with Explanation $F(1,135)=3.55$ $0.0618$ $-0.19$ $0.100$
Expl $F(1,115)=2.09$ $0.1511$ $0.04$ $0.031$
Disclosure $F(1,49)=4.71$ $0.0349$ $0.10$ $0.044$
protected Disclosure with Explanation $F(1,66)=0.23$ $0.6366$ $0.02$ $0.033$
Expl $F(1,227)=2.97$ $0.0863$ $0.02$ $0.018$
Disclosure $F(1,91)=1.73$ $0.1917$ $0.03$ $0.026$
Accuracy proxy Disclosure with Explanation $F(1,136)=2.22$ $0.1382$ $0.03$ $0.019$
Expl $F(1,115)=0.06$ $0.8126$ $0.01$ $0.096$
Disclosure $F(1,49)=0.01$ $0.9113$ $-0.01$ $0.125$
protected Disclosure with Explanation $F(1,66)=0.53$ $0.4674$ $0.08$ $0.104$
Expl $F(1,227)=8.8$ $0.0033$** $0.14$ $0.054$
Disclosure $F(1,91)=0.6$ $0.4400$ $0.06$ $0.072$
FNR proxy Disclosure with Explanation $F(1,136)=9.69$ $0.0023$** $0.20$ $0.063$
Expl $F(1,115)=0.7$ $0.4051$ $-0.08$ $0.107$
Disclosure $F(1,49)=0.73$ $0.3983$ $-0.12$ $0.144$
protected Disclosure with Explanation $F(1,66)=0.59$ $0.4458$ $-0.10$ $0.125$
Expl $F(1,227)=9.44$ $0.0024$** $-0.16$ $0.059$
Disclosure $F(1,91)=2.07$ $0.1533$ $-0.12$ $0.085$
FPR proxy Disclosure with Explanation $F(1,136)=8.79$ $0.0036$** $-0.22$ $0.073$
Table 2. Effects of dispositional trust in AI on different outcome metrics (Perception, Parity, Accuracy, FPR, and FNR).
@@
We consistently find that dispositional trust has no significant impact on fairness ratings in the protected conditions, while it significantly increases fairness ratings in the proxy conditions. In other words, when biases are direct, people are equally able to notice model biases even when they tend to trust AI in general; however, when biases are indirect, people with higher dispositional trust in AI are less likely to believe that the model is unfair. For participants’ fairness saliency, we see that there is never a significant effect in the case of direct bias, but higher dispositional trust significantly decreased the rate of disagreement only when considering disclosures without explanations.
We also find that, under models that consider the effect of explanations and disclosures with explanations, increases dispositional trust in AI significantly increased FPR and decreased FNR in the proxy conditions only. This is likely due to participants with higher trust in AI being more influenced by subtle indirect biases leading to lower acceptance rates for female applicants. This is also supported by the models measuring the effect of explanations on gender parity. Here, we see that an increased dispositional trust in AI significantly decreased parity under both types of bias.
Overall, people with a greater dispositional trust in AI tended to make more unfair decisions (when working with a biased model) and were less likely to notice indirect bias.
@@ B.2. Do explanations and disclosures affect learned trust?
In this section, we discuss the effect of our interventions—explanations, disclosures without explanations, and disclosures with explanations (Figure 1)—on learned trust over dispositional trust in AI generally. We perform statistical tests similar to the ones described in §5.2. We consider the different trust measures as the dependent variable and the treatment as the fixed effect term. We also control for the dispositional trust level as a fixed effect.
As seen in Table 6, we find that our treatments generally have no effect on trust ratings in models exhibiting direct bias, except for explanations alone leading to significantly lowered feelings that the AI system works well. In the case of indirect bias, we often see that full disclosure with explanations (and sometimes also full disclosure without explanations or bias disclosure with explanations) has a significant effect on learned trust. These effects demonstrate lowered feelings that the AI system works well as well as decreased feelings that the AI system can perform as well as an untrained human, decreased confidence in the system, decreased feelings of safety when relying on the system, and increased wariness of the AI system.
We also find that explanations alone significantly increase participant’s perception of model predictability in the proxy conditions but not the protected conditions. Without explanations, full bias and correlation disclosure also significantly increased predictability. With explanations, however, disclosures do not increase predictability, likely due to high predictability ratings even with explanations alone. This is to say that our models are already seen as relatively predictable when biases are direct, but when biases are indirect, explanations or disclosure of model bias and the model’s usage of the university feature help make the model more predictable.
better confidence predictable safe wary works well
Feature Phase Expl? Disclosure sig coef sig coef sig coef sig coef sig coef sig coef
$1$** $-0.61$** $-0.43$** $-0.51$** $-0.29$ $0.04$** $-0.43$
$1$** $-0.93$** $-0.66$** $-0.37$ $-0.26$ $0.10$** $-0.81$
$2$ Bias *** $-0.33$** $-0.23$** $-0.25$** $-0.20$ $0.02$** $-0.28$
Protected $2$ Bias *** $-0.51$** $-0.38$** $-0.19$** $-0.22$ $0.12$** $-0.48$
$1$** $-0.55$** $-0.34$** $-0.68$** $-0.22$ $-0.15$** $-0.32$
$1$** $-0.54$ $-0.19$ $0.01$ $0.03$ $-0.14$** $-0.29$
$2$ Bias *** $-0.39$** $-0.21$** $-0.29$ $-0.05$ $0.02$** $-0.25$
$2$ Full *** $-0.31$** $-0.28$** $-0.23$ $-0.13$ $0.01$** $-0.37$
$2$ Bias *** $-0.30$** $-0.29$ $0.02$ $-0.09$ $0.06$** $-0.30$
Proxy $2$ Full *** $-0.36$** $-0.20$ $-0.04$ $-0.08$ $0.00$** $-0.26$
Table 3. Comparison of dispositional trust vs learned trust in varied conditions and phases.
@@ B.3. Does dispositional trust differ from learned trust?
In the previous section, we discussed how interventions affected learned trust when controlling for baseline dispositional trust. Here, we study whether there is a significant difference between dispositional trust and learned trust in the biased models across the questions described in §4.1. These results are shown in Table 3.
We find that participants usually thought our model worked worse than AI does generally, that it inspired less confidence, and was less predictable. Participants also regarded our AI system as less safe than AI in general, but this is primarily true only in the case of direct bias. Surprisingly, participants did not consider our AI systems to be less safe than general in phase 1 when they were given explanations (which would have directly indicated that the system used gender as a feature to determine loan outcomes). Participants’ wariness of the biased models was not significantly different from their baseline wariness in AI.